# Bounded Hyperbolic Tangent: A Stable and Efficient Alternative to Pre-Layer Normalization in Large Language Models

### v1

@@ 4.5 Analysis on Scalability and Generalization
To validate BHyT in a more realistic pretraining setting, we extended training to 20B tokens and compared BHyT against Peri-LN. Table 4 reports results for both pretraining-only and post-SFT stages. BHyT achieves lower pretraining loss and perplexity, indicating that our method effectively suppresses depth-wise drift even with larger budgets. Additional results for the 1B model trained on 20B tokens can be found in Appendix D.5.
@@ Appendix B Proofs of the theorems
where $z_{t}$ is the activation of the $t$-th token after BHyT, and we let $W=W_{V}W_{O}$. Since the inputs and weights are zero-mean, the output $h_{\text{Attn}}$ also has a zero mean. Therefore, the variance of the $k$-th coordinate is $\mathrm{Var}((h_{\text{Attn}})_{k})=\mathbb{E}[(h_{\text{Attn}})_{k}^{2}]$. Substituting this into the definition of $\tilde{s}_{h_{\text{Attn}}}^{2}$, we can express the average variance using the trace of the covariance matrix:
@@ B.3 Approximation of the variance of the MLP layer output
Under the assumption that the activation function $\phi$ essentially acts as a scaling factor $\tau$ on the covariance matrix (which holds exactly for linear activations or as a linear approximation for non-linear ones), we substitute $\text{Cov}(v)\approx\tau\text{Cov}(u)$:
@@
While Llama architectures employ SwiGLU (shazeer2020glu) variants involving three matrices (gate, up, down), for theoretical tractability, we analyze the standard MLP formulation parameterized by $W_{1}$ and $W_{2}$. We assume this variance propagation logic generalizes to Gated units by treating the gating mechanism as part of the effective activation scaling factor $\tau$.
@@ B.4.3 Specialization to RMSNorm, LNS and BHyT
In this section, we instantiate the generic variance recursion from the assumptions and results of Theorems 3.4 and B.1, for three concrete choices of $f(\cdot)$: RMSNorm, LNS (sun2025curse), and BHyT. We follow the derivation in Appendix B.4.2.
@@ B.4.4 Variance recursion and product form
We first determine the variance of the RMSNorm output. For a vector $x$ with zero mean, RMSNorm is defined as $\text{RMSNorm}(x)=\gamma\odot\frac{x}{\sqrt{\frac{1}{d}\sum_{j=1}^{d}x_{j}^{2}}}$. Squaring and taking the expectation, and utilizing the symmetry of the isotropic input distribution (where $\mathbb{E}[\frac{x_{i}^{2}}{\frac{1}{d}\sum x_{j}^{2}}]=1$), we obtain:
@@
Let $f(x)$ be the LNS (sun2025curse) operation, defined as RMSNorm followed by a depth-dependent scalar $1/\sqrt{\ell}$ at layer $\ell$. Using the same notation as in Lemma B.4, the normalized variance gains for LNS are given by:
@@ C.1 Hyperparameters for Pretraining
  • DyT: learnable tanh-scaling parameter initialized using recommended values from the DyT (zhu2025transformers): the learnable tanh-scaling parameter is initialized following the values recommended in the original paper (zhu2025transformers): for Llama-1B, we set the scalars to 1.0 (before Attention), 0.5 (before MLP), and 0.5 (last layer), while for Llama-3B we use 0.2, 0.05, and 0.05, respectively.
@@ C.4 Hardware and system configuration
  • Pretrain and SFT framework: Llama-Factory (zheng2024llamafactory)
@@
  • Parallel computation: For efficient parallel computation, we use FlashAttention-2 (dao2022flashattention)
@@ D.1 Time to target loss analysis
RMSNorm 3.282 39.35 49K $\textit{31.28}_{\pm{0.32}}$ $62.52_{\pm{0.50}}$ $\textit{32.51}_{\pm{0.19}}$ $\textit{32.64}_{\pm{0.92}}$ $\textit{50.47}_{\pm{0.66}}$ $\textit{23.64}_{\pm{0.00}}$ $50.94_{\pm{0.50}}$ $40.57_{\pm{0.22}}$
Peri-LN 3.288 42.62 50K $\textbf{31.63}_{\pm{0.32}}$ $\textbf{63.07}_{\pm{0.43}}$ $32.05_{\pm{0.17}}$ $32.40_{\pm{0.80}}$ $49.41_{\pm{1.64}}$ $23.07_{\pm{0.00}}$ $\textit{58.05}_{\pm{0.34}}$ $\textit{41.38}_{\pm{0.12}}$
LNS 3.281 37.65 49K $31.02_{\pm{0.41}}$ $62.50_{\pm{0.48}}$ $\textbf{32.82}_{\pm{0.13}}$ $\textbf{32.65}_{\pm{0.70}}$ $50.02_{\pm{1.19}}$ $23.27_{\pm{0.00}}$ $46.15_{\pm{0.61}}$ $39.78_{\pm{0.35}}$
BHyT 3.289 33.22 46K $30.87_{\pm{0.19}}$ $\textit{62.75}_{\pm{0.33}}$ $32.36_{\pm{0.15}}$ $32.48_{\pm{0.18}}$ $\textbf{51.70}_{\pm{1.31}}$ $\textbf{23.94}_{\pm{0.00}}$ $\textbf{59.23}_{\pm{0.63}}$ $\textbf{41.90}_{\pm{0.15}}$
@@
RMSNorm 3.203 76.55 60K $31.01_{\pm{0.21}}$ $\textbf{66.57}_{\pm{0.41}}$ $\textbf{41.52}_{\pm{0.25}}$ $\textbf{33.92}_{\pm{0.33}}$ $50.50_{\pm{0.78}}$ $22.98_{\pm{0.00}}$ $37.94_{\pm{0.12}}$ $40.63_{\pm{0.18}}$
Peri-LN 3.165 88.72 60K $31.82_{\pm{0.38}}$ $64.52_{\pm{0.24}}$ $36.05_{\pm{0.11}}$ $32.28_{\pm{0.59}}$ $49.30_{\pm{0.52}}$ $22.99_{\pm{0.00}}$ $\textbf{57.46}_{\pm{0.31}}$ $\textit{42.06}_{\pm{0.25}}$
LNS 3.160 78.55 60K $\textit{31.88}_{\pm{0.42}}$ $64.68_{\pm{0.46}}$ $36.25_{\pm{0.10}}$ $\textit{32.45}_{\pm{0.53}}$ $\textit{51.18}_{\pm{0.75}}$ $\textbf{23.59}_{\pm{0.00}}$ $37.83_{\pm{0.08}}$ $39.69_{\pm{0.11}}$
BHyT 3.135 64.30 56K $\textbf{32.44}_{\pm{0.17}}$ $\textit{64.83}_{\pm{0.18}}$ $\textit{36.31}_{\pm{0.09}}$ $31.96_{\pm{0.36}}$ $\textbf{51.27}_{\pm{1.51}}$ $\textit{23.39}_{\pm{0.00}}$ $\textit{56.06}_{\pm{0.71}}$ $\textbf{42.32}_{\pm{0.31}}$
@@ D.4 Token generation speed
D.5 Pretraining results of Llama-1B on 20B tokens

### v2

@@ 4.5 Analysis on Scalability and Generalization
To validate BHyT in a more realistic pretraining setting, we extended training to 20B tokens and compared BHyT against Peri-LN. Table 4 reports results for both pretraining-only and post-SFT stages. BHyT achieves lower pretraining loss and perplexity, indicating that our method effectively suppresses depth-wise drift even with larger budgets. Additional results for the 1B model trained on 20B tokens can be found in Appendix E.1.
@@ Appendix B Proofs of the theorems
where $z_{t}$ is the activation of the $t$-th token after BHyT, and we let $W=W_{V}W_{O}$. Since the inputs and weights are zero-mean, the output $h_{\text{Attn}}$ also has a zero-mean. Therefore, the variance of the $k$-th coordinate is $\mathrm{Var}((h_{\text{Attn}})_{k})=\mathbb{E}[(h_{\text{Attn}})_{k}^{2}]$. Substituting this into the definition of $\tilde{s}_{h_{\text{Attn}}}^{2}$, we can express the average variance using the trace of the covariance matrix:
@@ B.3 Approximation of the variance of the MLP layer output
Under the assumption that the activation function $\phi$ essentially acts as a scaling factor $\tau$ on the covariance matrix (which holds exactly for linear activations or as a linear approximation for nonlinear ones), we substitute $\text{Cov}(v)\approx\tau\text{Cov}(u)$:
@@
While Llama architectures employ SwiGLU (Shazeer, 2020) variants involving three matrices (gate, up, down), for theoretical tractability, we analyze the standard MLP formulation parameterized by $W_{1}$ and $W_{2}$. We assume this variance propagation logic generalizes to Gated units by treating the gating mechanism as part of the effective activation scaling factor $\tau$.
@@ B.4.3 Specialization to RMSNorm, LNS and BHyT
In this section, we instantiate the generic variance recursion from the assumptions and results of Theorems 3.4 and B.1, for three concrete choices of $f(\cdot)$: RMSNorm, LNS (Sun et al., 2025), and BHyT. We follow the derivation in Appendix B.4.2.
@@ B.4.4 Variance recursion and product form
We first determine the variance of the RMSNorm output. For a vector $x$ with zero-mean, RMSNorm is defined as $\text{RMSNorm}(x)=\gamma\odot\frac{x}{\sqrt{\frac{1}{d}\sum_{j=1}^{d}x_{j}^{2}}}$. Squaring and taking the expectation, and utilizing the symmetry of the isotropic input distribution (where $\mathbb{E}[\frac{x_{i}^{2}}{\frac{1}{d}\sum x_{j}^{2}}]=1$), we obtain:
@@
Let $f(x)$ be the LNS (Sun et al., 2025) operation, defined as RMSNorm followed by a depth-dependent scalar $1/\sqrt{\ell}$ at layer $\ell$. Using the same notation as in Lemma B.4, the normalized variance gains for LNS are given by:
@@ C.1 Hyperparameters for Pretraining
  • DyT: learnable tanh-scaling parameter initialized using recommended values from the DyT (Zhu et al., 2025): the learnable tanh-scaling parameter is initialized following the values recommended in the original paper (Zhu et al., 2025): for Llama-1B, we set the scalars to 1.0 (before Attention), 0.5 (before MLP), and 0.5 (last layer), while for Llama-3B we use 0.2, 0.05, and 0.05, respectively.
@@ C.4 Hardware and system configuration
  • Pretrain and SFT framework: Llama-Factory (Zheng et al., 2024)
@@
  • Parallel computation: For efficient parallel computation, we use FlashAttention-2 (Dao, 2023)
@@ D.1 Time to target loss analysis
RMSNorm 3.282 39.35 49K $30.97_{\pm{0.20}}$ $\textit{62.89}_{\pm{0.74}}$ $\textit{32.77}_{\pm{0.09}}$ $32.32_{\pm{0.94}}$ $\textbf{50.54}_{\pm{0.48}}$ $\textbf{25.70}_{\pm{0.01}}$ $56.26_{\pm{0.44}}$ $41.64_{\pm{0.10}}$
Peri-LN 3.288 42.62 50K $\textbf{31.63}_{\pm{0.32}}$ $\textbf{63.07}_{\pm{0.43}}$ $32.05_{\pm{0.17}}$ $32.40_{\pm{0.80}}$ $49.41_{\pm{1.64}}$ $24.91_{\pm{0.00}}$ $\textit{58.05}_{\pm{0.34}}$ $41.65_{\pm{0.12}}$
LNS 3.281 37.65 49K $\textit{31.39}_{\pm{0.24}}$ $62.94_{\pm{0.64}}$ $\textbf{32.80}_{\pm{0.16}}$ $\textit{33.50}_{\pm{0.58}}$ $\textit{50.37}_{\pm{0.91}}$ $24.88_{\pm{0.01}}$ $56.67_{\pm{0.37}}$ $\textit{41.79}_{\pm{0.26}}$
[…]
@@
RMSNorm 3.203 76.55 60K $31.01_{\pm{0.21}}$ $\textbf{66.57}_{\pm{0.41}}$ $\textbf{41.52}_{\pm{0.25}}$ $\textbf{33.92}_{\pm{0.33}}$ $50.50_{\pm{0.78}}$ $25.67_{\pm{0.02}}$ $37.94_{\pm{0.12}}$ $40.89_{\pm{0.18}}$
Peri-LN 3.165 88.72 60K $31.82_{\pm{0.38}}$ $64.52_{\pm{0.24}}$ $36.05_{\pm{0.11}}$ $32.28_{\pm{0.59}}$ $49.30_{\pm{0.52}}$ $25.99_{\pm{0.00}}$ $\textit{57.46}_{\pm{0.31}}$ $\textit{42.49}_{\pm{0.25}}$
LNS 3.160 78.55 60K $\textbf{31.88}_{\pm{0.42}}$ $64.68_{\pm{0.46}}$ $36.25_{\pm{0.10}}$ $\textit{32.45}_{\pm{0.53}}$ $\textit{51.18}_{\pm{0.75}}$ $\textbf{26.87}_{\pm{0.01}}$ $37.83_{\pm{0.08}}$ $41.16_{\pm{0.11}}$
BHyT 3.135 64.30 56K $\textit{31.84}_{\pm{0.15}}$ $\textit{65.08}_{\pm{0.39}}$ $\textit{36.48}_{\pm{0.04}}$ $31.76_{\pm{0.33}}$ $\textbf{51.62}_{\pm{0.66}}$ $25.70_{\pm{0.01}}$ $\textbf{60.84}_{\pm{0.29}}$ $\textbf{43.33}_{\pm{0.16}}$
@@ D.4 Token generation speed
Appendix E Pretraining and SFT results for Llama-1B
In the main text, we focused on the scalability of BHyT using the Llama-3B model. Here, we provide the detailed experimental results for the Llama-1B model under the same standard pretraining (Standard C4 budget) and SFT settings.
Table 9: Pretraining (PT)-only evaluation for Llama-1B. BHyT consistently achieves the lowest pretraining loss and superior average accuracy compared to baselines.
Llama-1B
Method
PT Train
Loss
PT Eval
Loss
PT Eval
PPL
Arc-e PIQA Hellaswag OpenBookQA Winogrande MMLU BoolQ Avg.
[…]
Table 10: Supervised fine-tuning (SFT) results for Llama-1B. BHyT maintains stable optimization and attains competitive accuracy on downstream benchmarks.
Llama-1B
Method
PT Eval
Loss
SFT Train
Loss
SFT Eval
Loss
Arc-e PIQA Hellaswag OpenBookQA Winogrande MMLU BoolQ Avg.
Regarding the inference speed, detailed comparisons for Llama-1B across different generation lengths are provided in Appendix D.4 and Table 8.
E.1 Pretraining results of Llama-1B on 20B tokens